# WENDY: Gene Regulatory Network Inference with Covariance Dynamics

### v1

@@ 1 Introduction
In general, a gene is transcribed into mRNA and then translated into proteins. This process, known as gene expression, commonly employs mRNA count or protein count to denote the expression level. In addition to directly changing cell phenotypes [66, 17], influencing extracellular processes [5], or even manipulating macroscopic neurological circuitry [47, 79], certain proteins can affect the transcription of other genes (mutual regulation) or their own corresponding genes (autoregulation). Genes and their regulatory relationships form a gene regulatory network (GRN).
@@
Determining the GRN structure is a central problem in biology, as it reveals how a living organism is maintained [4], and provides control of essential biological processes [86], especially treating cancer [54, 16]. However, directly establishing the GRN using traditional technologies is extremely difficult since they cannot measure the expression levels of many genes within the same cell. Instead, numerous methods have been developed to infer the GRN structure from gene expression data. Particularly, recent advancements in single-cell RNA-sequencing technologies have made it possible to profile the whole transcriptome of single cells at large-scale. However, single-cell RNA-seq can only measure one time point because cells have to be killed during the experimental process, making it challenging to study gene regulation relationships that require multiple observations over time.
@@
In this paper, we focus on a specific data type arising from the following setup: First, implement an intervention that affects gene expression (e.g., drugs). Then measure the expression level (generally mRNA count) of $n$ genes for different single cells at multiple time points, and select the data from time points where the expression has not yet reached a stationary state. Since gene expression at the single-cell level is stochastic, for each time point, we obtain many samples of an $n$-dimensional random vector. However, since we need to kill a cell before measuring its gene expression levels, one cell can only be measured once. Thus, we measure different cells at different time points, and we do not have a joint distribution for gene expression at different time points. Although this approach has become common in recent experimental research [13], and it provides more informative data compared to most other approaches, to our knowledge, there is only one inference method developed specifically for this data type, SINCERITIES [61]. A major limitation of SINCERITIES is that it requires data from at least six time points to perform well. Additionally, for single-cell expression data of $n$ genes over $T$ time points, this method only extracts $n(T-1)$ numbers for further analyses, implying low data utilization efficiency. There have been many inference methods for single-cell time series gene expression data, where the joint distribution of expression levels at different time points is known [37, 101]. Since obtaining the joint distribution of gene expression is difficult, such methods are usually not practically applicable. There are also many inference methods for single-cell gene expression data measured at a single time point [8, 95], or bulk level gene expression data measured at multiple time points after interventions [62, 51, 87]. These data types are more common because of their low cost. Nevertheless, they provide less information compared to the data type we examine in this study, namely, time series data from single-cell gene expression. Therefore, while it is feasible to convert our considered data type into these more common forms and use corresponding inference methods, such transformations result in a significant loss of the rich information inherent in the original dataset.
@@ 2.1 A framework for data classification
There are different types of gene expression data that can be used to infer the GRN structure. Different data types correspond to different inference methods. We first present a framework for classifying related data types, modified from the framework by Wang and Wang [85]. See Table 1 for this classification framework. There are different dimensions to classify data types.
@@
In practice, repeating the same bulk level measurement can still lead to different values, making some researchers regard such data as stochastic and apply inference methods designed for single-cell data [8]. Nevertheless, at bulk level, randomness from single cells is averaged out, and the different values from bulk level measurement can only come from systematic differences, such as different cell phenotypes or different environmental factors. Such unobserved systematic differences can affect multiple genes and make them correlated, although these genes might not have direct regulatory relations. Therefore, we do not consider bulk level data that have different values for the same measurement.
@@
In practice, measuring the expression levels of many genes is destructive, and we cannot measure the same cell more than once. If we only want to measure the expression level of a single gene, there are some techniques (fluorescent proteins [91], etc.) that can measure the same cell multiple times. Another approach is to measure the amount of spliced and unspliced mRNAs, which provides both the current expression level and an approximation of its time derivative (RNA velocity [44]). This approach provides two measurements of the same cell, and some inference methods for time series data can be applied.
@@
Table 1: Classification of data types regarding GRN inference, modified from the framework by Wang and Wang [85]. The data types are classified by different dimensions: (1) The gene expression is at stationary, or is driven away from stationary by an intervention (on general genes that we cannot choose, or specific genes that we can choose); (2) Measure at one time point or multiple time points (time series); (3) Measure the average over many genes (bulk level) or on single-cell level. When measuring at multiple time points on single-cell level, one more dimension is whether we have the joint distribution over different time points. For different data types (scenarios), we study whether the GRN structure can be inferred. For each data type, No means that the GRN structure cannot be inferred. Ditto means that there are no specific inference methods, but the GRN structure can be inferred using the same method for a less informative data type. Yes means specific inference methods exist. Scenario 11 only has an inference method that also requires the data in Scenario 1. We focus on Scenario 9, which is not as well-studied and has only one specific inference method.
@@ 2.2 Known inference methods for different data types
For bulk level data types, since we only have a single deterministic value for each gene, it is difficult to obtain the correlation between genes. For Scenarios 1, 3, 6, the GRN structure cannot be inferred. For Scenarios 8 and 13, we can regard the gene expression time series as solution trajectories of an ordinary differential equation (ODE) system. If we assume that the ODE system is linear [62] or has certain nonlinear forms [51], we can discretize the ODE system into an algebraic equation system and use regression to infer the ODE parameters. Here the ODE parameters represent the GRN. We can infer all the edges, including the directions. For Scenario 11, one can add an intervention on each gene and observe which genes (descendants of this gene in the GRN) are also affected. Such ancestor-descendant relationships can be used to partially infer the GRN structure [85]. Not all regulatory relationships (edges) can be inferred, but one can infer at least $n-1$ edges for a GRN with $n$ genes.
@@
For single-cell level one-time data types (Scenarios 2, 7, 12), there have been numerous GRN inference methods for Scenario 2, while Scenarios 7 and 12 generally do not have extra information that supports specific inference methods. For Scenario 2, most inference methods turn the GRN inference problem into a feature selection problem: select genes whose levels can be best used to predict the level of the target gene. Then such selected genes might have regulatory relationships with the target gene. The selection can be made by calculating certain quantities between the target gene and the candidate genes that measure their similarity: covariance [60], mutual information [8], or other information theory quantities [95]. Besides, one can apply regression [33], decision tree [38], or other machine learning and deep learning [70, 100, 103] methods to directly select out genes that can predict the target gene. Regularization terms (e.g., $L_{1}$ [32] and $L_{2}$ [41] regularizers) can be added to the regression to make the result sparse. Besides the idea of feature selection, another approach is to build probabilistic models (Bayesian network [2, 45], stochastic differential equation (SDE) [82], and others [49, 12]), and use likelihood to determine the most probable network. Since the number of candidate networks is large, a common solution is to apply Markov chain Monte Carlo (MCMC) to approximate the network probabilities [55, 2, 1, 31, 102]. Deterministic models, such as Boolean networks [48], can also be used. There is a well-developed platform that combines different inference methods for Scenario 2 [89]. One problem of Scenario 2 is to determine the direction of a regulatory relationship, since if the level of $V_{i}$ can predict the level of $V_{j}$, then generally the level of $V_{j}$ can also predict the level of $V_{i}$. To solve this problem, one can add specific interventions (Scenario 12) on $V_{i}$ to see whether $V_{j}$ is affected.
@@
For single-cell level time series data types without joint distribution of different time points (Scenarios 4, 9, 14), it is common to treat Scenario 4 as Scenario 2 (treat data at different time points separately), and treat Scenarios 9, 14 as Scenario 8 (average over different cells), and apply corresponding inference methods. The only GRN inference method we know that works specifically for Scenario 9 (but not any less informative scenarios) is SINCERITIES [61], which considers the Kolmogorov–Smirnov distance between the distributions of the same gene at two time points, and then applies linear regression, similar to methods for Scenario 8. There are two other regression-based methods, Harissa and CARDAMOM [35, 77], that essentially work for Scenario 2, but can partially (and insufficiently) integrate the time information when work for Scenario 9.
@@
For single-cell level time series data types with joint distribution of different time points (Scenarios 5, 10, 15), there have been numerous GRN inference methods for Scenario 5, while Scenarios 10 and 15 generally do not have extra information that support specific inference methods. For Scenario 5, most inference methods are similar to those for Scenario 2, especially those methods on regression [96], tree-based feature selection [37, 101], or more advanced machine learning tools [57, 3], although there are also inference methods based on more complicated biological models [39]. One common approach is to model the gene expression by a vector autoregressive model (either linear or nonlinear) [27, 28, 71], and then use Granger causality to determine whether one gene directly regulates another gene [29, 30, 56, 98]. This approach solves the famous “correlation does not imply causation” problem from two aspects. First, when gene $V_{i}$ and gene $V_{j}$ are correlated, it is possible that they are not directly regulating each other, such as $V_{i}\leftarrow V_{k}\to V_{j}$ or $V_{i}\to V_{k}\to V_{j}$. In this case, given the values of $V_{k}$, $V_{i}$ cannot provide more information for $V_{j}$, and Granger causality can determine that $V_{i}$ does not directly regulate $V_{j}$. Second, when there is directly regulation between $V_{i}$ and $V_{j}$, we do not know whether $V_{i}$ is the cause or the result of $V_{j}$. Since Granger causality determines whether the past of $V_{i}$ contains unique information of the future of $V_{j}$, the direction of regulation is also known, since causality can only travel forward along time. This explains why most methods for Scenario 5 can determine the direction of regulations, different from their analogies for Scenario 2.
@@
Besides treating a more informative data type as a less informative data type, one can also use certain methods to transform a less informative data type into a more informative data type, provided there are certain assumptions about the underlying systems. For instance, from single-cell one-time data (Scenarios 2, 7, 12), one can construct the pseudotime to transform the data into time series data [67, 74], and apply corresponding inference methods [22].
@@
Different methods need different assumptions regarding gene expression and gene regulation. For instance, some methods need the gene expression dynamics to be linear [62], and some other methods need the GRN to have no directed cycle [95]. Besides, different methods have different inference abilities: some methods can determine all edges, including the direction [37], while some other methods can only partially determine some edges, and/or cannot determine the edge direction [85].
@@
Most GRN inference methods can only determine regulations between genes, but not autoregulation. Autoregulation inference methods generally need stronger model assumptions [92], more informative data types [25], or only produce partial results [83].
@@
The above discussion only considers the situation of inferring GRN after obtaining all data. Another situation is to design intervention experiments, so that the GRN can be inferred with the minimal cost [18].
@@
Readers may refer to other reviews for more details about GRN inference methods for different scenarios [7, 65, 99, 59, 85] or for other data types (besides mRNA/protein count) that can help with GRN inference, such as ChIP-seq and ATAC-seq [24, 6].
@@ 3.1 Mathematical model of gene expression
Here $c_{1,i}$ means the basal synthesis rate of $X_{i}$, and $c_{2,i}X_{i}$ means the total degradation rate of $X_{i}$. The function $f$ that reflects the interaction between genes can take any form with any number of unknown parameters. The noise term is generally Gaussian. We think that the random fluctuation is proportional to $X_{i}$, and decide to follow Pinna et al. and Papili Gao et al. [64, 61] to set the noise term to be $\sigma_{i}X_{i}\mathrm{d}W_{i}(t)$, where $\sigma_{i}$ is an unknown constant, and $W_{i}(t)$ is a standard Brownian motion.
Since $f$ in Eq. 1 is an arbitrary function, it is impossible to determine its parameters. We need to add restrictions on the form of $f$. Besides single gene regulations $V_{j}\to V_{i}$, transcription factors can bind to enhancers, whose signals are transmitted by coactivators (e.g., mediators) to promoters, which can bind to general transcription factors and regulate the expression of certain genes [72, 40, 11]. Therefore, to model the complicated gene expression, at least we need to consider cooperative regulations with two genes, $V_{j}+V_{k}\to V_{i}$. This means that Eq. 1 becomes
@@
To determine the GRN, we need to determine all coefficients $a_{j\to i},b_{j,k\to i}$, whose total number is $n^{3}/2+n^{2}/2$ for $n$ genes. For single-cell gene expression data, at each time point, the data set is a matrix of $m$ cells $\times$ $n$ genes. From the data, we can estimate the parameters of the joint distribution of $n$ genes. If this joint distribution is Gaussian, we can obtain $n$ parameters for the mean, and $n(n+1)/2$ parameters for the (co)variance (since the covariance matrix is symmetric), which fully determine the distribution. Even though the joint distribution of $n$ genes might not be Gaussian in reality, the number of parameters that can be estimated from expression data at each time point should be at the level of $n^{2}/2$. Therefore, to determine $n^{3}/2+n^{2}/2$ GRN parameters in Eq. 2, we need at least $n$ time points. This generally does not hold in reality, since a small GRN of interest might contain tens of genes, but the number of time points for type 9 data is commonly only a few [34, 76, 19]. Thus we have to further (over)simplify Eq. 2 and drop cooperative terms:
@@
Gene expression has many biochemical subtleties, and Eq. 7 is certainly an oversimplification of Eq. 1. Here the simplification from Eq. 1 to Eq. 3 is inevitable, since type 9 data generally do not have many different time points, and cannot provide enough information to solve a model with too many unknown parameters (as in Eq. 2). If we want to learn some knowledge about the GRN from type 9 data, some simplification is necessary. The simplifications from Eq. 3 to Eq. 7 are only for numerical purposes. In Subsection 4.3, We will see that although our method is derived from Eq. 7, it has good performance on data generated by Eq. 11, which is in the form of Eq. 3. Therefore, the simplification from Eq. 3 to Eq. 7 does not harm the generalizability of our method. Besides, although our method cannot determine autoregulation $V_{i}\to V_{i}$, it can determine feedback loop $V_{i}\to V_{j}\to V_{i}$, which is the fundamental mechanism of some “autoregulations” observed in experiments [20].

### v2

# WENDY: Covariance Dynamics Based Gene Regulatory Network Inference

@@ 1 Introduction
In general, a gene is transcribed into mRNA and then translated into proteins. This process, known as gene expression, commonly employs mRNA count or protein count to denote the expression level. In addition to directly changing cell phenotypes (Qian and Cheng,, 2020; Cheng et al., 2023b, ), influencing extracellular processes (Axelrod et al.,, 2023), or even manipulating macroscopic neurological circuitry (Li et al.,, 2021; Vijayan et al.,, 2022), certain proteins can affect the transcription of other genes (mutual regulation) or their own corresponding genes (autoregulation). Genes and their regulatory relationships form a gene regulatory network (GRN).
@@
Determining the GRN structure is a central problem in biology, as it reveals how a living organism is maintained (Axelrod et al.,, 2021), and provides control of essential biological processes (Wang,, 2020), especially treating cancer (McDonald et al.,, 2023; Cheng et al., 2023a, ). However, directly establishing the GRN using traditional technologies is extremely difficult since they cannot measure the expression levels of many genes within the same cell. Instead, numerous methods have been developed to infer the GRN structure from gene expression data. Particularly, recent advancements in single-cell RNA-sequencing technologies have made it possible to profile the whole transcriptome of single cells at large-scale. However, single-cell RNA-seq can only measure one time point because cells have to be killed during the experimental process, making it challenging to study gene regulation relationships that require multiple observations over time.
@@
In this paper, we focus on a specific data type arising from the following setup: First, implement an intervention that affects gene expression (e.g., drugs). Then measure the expression level (generally mRNA count) of $n$ genes for different single cells at multiple time points, and select the data from time points where the expression has not yet reached a stationary state. Since gene expression at the single-cell level is stochastic, for each time point, we obtain many samples of an $n$-dimensional random vector. However, since we need to kill a cell before measuring its gene expression levels, one cell can only be measured once. Thus, we measure different cells at different time points, and we do not have a joint distribution for gene expression at different time points. Although this approach has become common in recent experimental research (Chakraborty et al.,, 2021), and it provides more informative data compared to most other approaches, to our knowledge, there is only one inference method developed specifically for this data type, SINCERITIES (Papili Gao et al.,, 2018). A major limitation of SINCERITIES is that it requires data from at least six time points to perform well. Additionally, for single-cell expression data of $n$ genes over $T$ time points, this method only extracts $n(T-1)$ numbers for further analyses, implying low data utilization efficiency. There have been many inference methods for single-cell time series gene expression data, where the joint distribution of expression levels at different time points is known (Huynh-Thu and Geurts,, 2018; Zheng et al.,, 2019). Since obtaining the joint distribution of gene expression is difficult, such methods are usually not practically applicable. There are also many inference methods for single-cell gene expression data measured at a single time point (Basso et al.,, 2005; Zhang et al.,, 2012), or bulk level gene expression data measured at multiple time points after interventions (Perrin et al.,, 2003; Ma et al.,, 2020; Wang et al.,, 2022). These data types are more common because of their low cost. Nevertheless, they provide less information compared to the data type we examine in this study, namely, time series data from single-cell gene expression. Therefore, while it is feasible to convert our considered data type into these more common forms and use corresponding inference methods, such transformations result in a significant loss of the rich information inherent in the original dataset.
@@ 2.1 A framework for data classification
There are different types of gene expression data that can be used to infer the GRN structure. Different data types correspond to different inference methods. We first present a framework for classifying related data types, modified from the framework by Wang and Wang, (2022). See Table 1 for this classification framework. There are different dimensions to classify data types.
@@
In practice, repeating the same bulk level measurement can still lead to different values, making some researchers regard such data as stochastic and apply inference methods designed for single-cell data (Basso et al.,, 2005). Nevertheless, at bulk level, randomness from single cells is averaged out, and the different values from bulk level measurement can only come from systematic differences, such as different cell phenotypes or different environmental factors. Such unobserved systematic differences can affect multiple genes and make them correlated, although these genes might not have direct regulatory relations. Therefore, we do not consider bulk level data that have different values for the same measurement.
@@
In practice, measuring the expression levels of many genes is destructive, and we cannot measure the same cell more than once. If we only want to measure the expression level of a single gene, there are some techniques (fluorescent proteins (Wu et al.,, 2011), etc.) that can measure the same cell multiple times. Another approach is to measure the amount of spliced and unspliced mRNAs, which provides both the current expression level and an approximation of its time derivative (RNA velocity (La Manno et al.,, 2018)). This approach provides two measurements of the same cell, and some inference methods for time series data can be applied.
@@
Table 1: Classification of data types regarding GRN inference, modified from the framework by Wang and Wang, (2022). The data types are classified by different dimensions: (1) The gene expression is at stationary, or is driven away from stationary by an intervention (on general genes that we cannot choose, or specific genes that we can choose); (2) Measure at one time point or multiple time points (time series); (3) Measure the average over many genes (bulk level) or on single-cell level. When measuring at multiple time points on single-cell level, one more dimension is whether we have the joint distribution over different time points. For different data types (scenarios), we study whether the GRN structure can be inferred. For each data type, No means that the GRN structure cannot be inferred. Ditto means that there are no specific inference methods, but the GRN structure can be inferred using the same method for a less informative data type. Yes means specific inference methods exist. Scenario 11 only has an inference method that also requires the data in Scenario 1. We focus on Scenario 9, which is not as well-studied and has only one specific inference method.
@@ 2.2 Known inference methods for different data types
For bulk level data types, since we only have a single deterministic value for each gene, it is difficult to obtain the correlation between genes. For Scenarios 1, 3, 6, the GRN structure cannot be inferred. For Scenarios 8 and 13, we can regard the gene expression time series as solution trajectories of an ordinary differential equation (ODE) system. If we assume that the ODE system is linear (Perrin et al.,, 2003) or has certain nonlinear forms (Ma et al.,, 2020), we can discretize the ODE system into an algebraic equation system and use regression to infer the ODE parameters. Here the ODE parameters represent the GRN. We can infer all the edges, including the directions. For Scenario 11, one can add an intervention on each gene and observe which genes (descendants of this gene in the GRN) are also affected. Such ancestor-descendant relationships can be used to partially infer the GRN structure (Wang and Wang,, 2022). Not all regulatory relationships (edges) can be inferred, but one can infer at least $n-1$ edges for a GRN with $n$ genes.
@@
For single-cell level one-time data types (Scenarios 2, 7, 12), there have been numerous GRN inference methods for Scenario 2, while Scenarios 7 and 12 generally do not have extra information that supports specific inference methods. For Scenario 2, most inference methods turn the GRN inference problem into a feature selection problem: select genes whose levels can be best used to predict the level of the target gene. Then such selected genes might have regulatory relationships with the target gene. The selection can be made by calculating certain quantities between the target gene and the candidate genes that measure their similarity: covariance (Nouri et al.,, 2023), mutual information (Basso et al.,, 2005), or other information theory quantities (Zhang et al.,, 2012). Besides, one can apply regression (Haury et al.,, 2012), decision tree (Huynh-Thu et al.,, 2010), or other machine learning and deep learning (Shrivastava et al.,, 2022; Zhao et al.,, 2022; Zhong and Tibshirani,, 2019) methods to directly select out genes that can predict the target gene. Regularization terms (e.g., $L_{1}$ (Gustafsson et al.,, 2005) and $L_{2}$ (Kamimoto et al.,, 2023) regularizers) can be added to the regression to make the result sparse. Besides the idea of feature selection, another approach is to build probabilistic models (Bayesian network (Agostinho et al.,, 2015; Lee et al.,, 2019), stochastic differential equation (SDE) (Wang et al.,, 2023), and others (Liu et al.,, 2016; Burdziak et al.,, 2023)), and use likelihood to determine the most probable network. Since the number of candidate networks is large, a common solution is to apply Markov chain Monte Carlo (MCMC) to approximate the network probabilities (Morrissey et al.,, 2010; Agostinho et al.,, 2015; Aalto et al.,, 2020; Ghosh and Zhong,, 2021; Zhong,, 2021). Deterministic models, such as Boolean networks (Lim et al.,, 2016), can also be used. There is a well-developed platform that combines different inference methods for Scenario 2 (Wen et al.,, 2023). One problem of Scenario 2 is to determine the direction of a regulatory relationship, since if the level of $V_{i}$ can predict the level of $V_{j}$, then generally the level of $V_{j}$ can also predict the level of $V_{i}$. To solve this problem, one can add specific interventions (Scenario 12) on $V_{i}$ to see whether $V_{j}$ is affected.
@@
For single-cell level time series data types without joint distribution of different time points (Scenarios 4, 9, 14), it is common to treat Scenario 4 as Scenario 2 (treat data at different time points separately), and treat Scenarios 9, 14 as Scenario 8 (average over different cells), and apply corresponding inference methods. The only GRN inference method we know that works specifically for Scenario 9 (but not any less informative scenarios) is SINCERITIES (Papili Gao et al.,, 2018), which considers the Kolmogorov–Smirnov distance between the distributions of the same gene at two time points, and then applies linear regression, similar to methods for Scenario 8. There are two other regression-based methods, Harissa and CARDAMOM (Herbach,, 2023; Ventre,, 2021), that essentially work for Scenario 2, but can partially (and insufficiently) integrate the time information when work for Scenario 9.
@@
For single-cell level time series data types with joint distribution of different time points (Scenarios 5, 10, 15), there have been numerous GRN inference methods for Scenario 5, while Scenarios 10 and 15 generally do not have extra information that support specific inference methods. For Scenario 5, most inference methods are similar to those for Scenario 2, especially those methods on regression (Zhang et al., 2021a, ), tree-based feature selection (Huynh-Thu and Geurts,, 2018; Zheng et al.,, 2019), or more advanced machine learning tools (Nauta et al.,, 2019; Atanackovic et al.,, 2023), although there are also inference methods based on more complicated biological models (Huynh-Thu and Sanguinetti,, 2015). One common approach is to model the gene expression by a vector autoregressive model (either linear or nonlinear) (Fujita et al., 2007a, ; Fujita et al., 2007b, ; Siggiridou and Kugiumtzis,, 2015), and then use Granger causality to determine whether one gene directly regulates another gene (Fujita et al., 2010a, ; Fujita et al., 2010b, ; Nagarajan and Upreti,, 2010; Zhang et al.,, 2010). This approach solves the famous “correlation does not imply causation” problem from two aspects. First, when gene $V_{i}$ and gene $V_{j}$ are correlated, it is possible that they are not directly regulating each other, such as $V_{i}\leftarrow V_{k}\to V_{j}$ or $V_{i}\to V_{k}\to V_{j}$. In this case, given the values of $V_{k}$, $V_{i}$ cannot provide more information for $V_{j}$, and Granger causality can determine that $V_{i}$ does not directly regulate $V_{j}$. Second, when there is directly regulation between $V_{i}$ and $V_{j}$, we do not know whether $V_{i}$ is the cause or the result of $V_{j}$. Since Granger causality determines whether the past of $V_{i}$ contains unique information of the future of $V_{j}$, the direction of regulation is also known, since causality can only travel forward along time. This explains why most methods for Scenario 5 can determine the direction of regulations, different from their analogies for Scenario 2.
@@
Besides treating a more informative data type as a less informative data type, one can also use certain methods to transform a less informative data type into a more informative data type, provided there are certain assumptions about the underlying systems. For instance, from single-cell one-time data (Scenarios 2, 7, 12), one can construct the pseudotime to transform the data into time series data (Reid and Wernisch,, 2016; Street et al.,, 2018), and apply corresponding inference methods (Deshpande et al.,, 2022).
@@
Different methods need different assumptions regarding gene expression and gene regulation. For instance, some methods need the gene expression dynamics to be linear (Perrin et al.,, 2003), and some other methods need the GRN to have no directed cycle (Zhang et al.,, 2012). Besides, different methods have different inference abilities: some methods can determine all edges, including the direction (Huynh-Thu and Geurts,, 2018), while some other methods can only partially determine some edges, and/or cannot determine the edge direction (Wang and Wang,, 2022).
@@
Most GRN inference methods can only determine regulations between genes, but not autoregulation. Autoregulation inference methods generally need stronger model assumptions (Xing and Van Der Laan,, 2005), more informative data types (Feigelman et al.,, 2016), or only produce partial results (Wang and He,, 2023).
@@
The above discussion only considers the situation of inferring GRN after obtaining all data. Another situation is to design intervention experiments, so that the GRN can be inferred with the minimal cost (Cho et al.,, 2016).
@@
Readers may refer to other reviews for more details about GRN inference methods for different scenarios (Barbosa et al.,, 2018; Pratapa et al.,, 2020; Zhao et al.,, 2021; Nguyen et al.,, 2021; Wang and Wang,, 2022) or for other data types (besides mRNA/protein count) that can help with GRN inference, such as ChIP-seq and ATAC-seq (Erbe et al.,, 2022; Badia-i Mompel et al.,, 2023).
@@ 3.1 Mathematical model of gene expression
Here $c_{1,i}$ means the basal synthesis rate of $X_{i}$, and $c_{2,i}X_{i}$ means the total degradation rate of $X_{i}$. The function $f$ that reflects the interaction between genes can take any form with any number of unknown parameters. The noise term is generally Gaussian. We think that the random fluctuation is proportional to $X_{i}$, and decide to follow Pinna et al., (2010); Papili Gao et al., (2018) to set the noise term to be $\sigma_{i}X_{i}\mathrm{d}W_{i}(t)$, where $\sigma_{i}$ is an unknown constant, and $W_{i}(t)$ is a standard Brownian motion.
Since $f$ in Eq. 1 is an arbitrary function, it is impossible to determine its parameters. We need to add restrictions on the form of $f$. Besides single gene regulations $V_{j}\to V_{i}$, transcription factors can bind to enhancers, whose signals are transmitted by coactivators (e.g., mediators) to promoters, which can recruit additional transcription factors and regulate the expression of one or many genes (Smith et al.,, 2023; Kamal et al.,, 2023). Regulation of gene expression by enhancers can occur over long distances, and in some instances, on different chromosomes. Some enhancers can regulate more than 1 gene (Bravo González-Blas et al.,, 2023). Therefore, to model the complicated gene expression, at least we need to consider cooperative regulations with two genes, $V_{j}+V_{k}\to V_{i}$. This means that Eq. 1 becomes
@@
To determine the GRN, we need to determine all coefficients $a_{j\to i},b_{j,k\to i}$, whose total number is $n^{3}/2+n^{2}/2$ for $n$ genes. For single-cell gene expression data, at each time point, the data set is a matrix of $m$ cells $\times$ $n$ genes. From the data, we can estimate the parameters of the joint distribution of $n$ genes. If this joint distribution is Gaussian, we can obtain $n$ parameters for the mean, and $n(n+1)/2$ parameters for the (co)variance (since the covariance matrix is symmetric), which fully determine the distribution. Even though the joint distribution of $n$ genes might not be Gaussian in reality, the number of parameters that can be estimated from expression data at each time point should be at the level of $n^{2}/2$. Therefore, to determine $n^{3}/2+n^{2}/2$ GRN parameters in Eq. 2, we need at least $n$ time points. This generally does not hold in reality, since a small GRN of interest might contain tens of genes, but the number of time points for type 9 data is commonly only a few (Hayashi et al.,, 2018; Treutlein et al.,, 2016; Chu et al.,, 2016). Thus we have to further (over)simplify Eq. 2 and drop cooperative terms:
@@
Gene expression has many biochemical subtleties, and Eq. 7 is certainly an oversimplification of Eq. 1. Here the simplification from Eq. 1 to Eq. 3 is inevitable, since type 9 data generally do not have many different time points, and cannot provide enough information to solve a model with too many unknown parameters (as in Eq. 2). If we want to learn some knowledge about the GRN from type 9 data, some simplification is necessary. The simplifications from Eq. 3 to Eq. 7 are only for numerical purposes. In Subsection 4.3, We will see that although our method is derived from Eq. 7, it has good performance on data generated by Eq. 11, which is in the form of Eq. 3. Therefore, the simplification from Eq. 3 to Eq. 7 does not harm the generalizability of our method. Besides, although our method cannot determine autoregulation $V_{i}\to V_{i}$, it can determine feedback loop $V_{i}\to V_{j}\to V_{i}$, which is the fundamental mechanism of some “autoregulations” observed in experiments (Crews and Pearson,, 2009).